What is the pH of a 1.13M solution of ascorbic acid? Ka1 = 8.0 x 10-5 Ka2 = 1.6 x 10-12
Answer #1

Setting up an ICE table by assuming an arbitrary Dibasic acid H2A that dissociates by a fraction \alpha at equilibrium:

                                                H2A      \rightarrow      HA-           +      H+

Initial concentration:              1.13 M                         0                            0

equilibrium concentration: 1.13(1 - \alpha) 1.13\alpha 1.13\alpha

The Ka1 value is equated to determine \alpha:

Ka1 = [HA-] * [H+]/[H2A]

8.0 x 10-5 = 1.13\alpha * 1.13\alpha / {1.13 * (1 - \alpha)}

Since \alpha << 1, ( 1 - \alpha) = 1

Solving, we get: \alpha = 8.41 * 10-3

Since \alpha is already so negligible and Ka1 is nearly 106 times Ka2, the dissociation of HA- can be neglected.

[H+] = 1.13 * \alpha = 1.13 * 8.41 * 10-3 = 9.50 * 10-3

pH = -log([H+]) = - log (9.50 * 10-3) = 2.022